# Parse Trees Guided LLM Prompt Compression

### v1

@@ VI-B Comparative Study: Under Various Compression Ratios
The compression ratio is a vital aspect in the prompt compression problem. To further investigate the effectiveness of these methods across a range of compression ratios, we carry out the experiments with nine different compression ratios, i.e., $\tau=0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9$. The results for BBCnews, truncate arXiv, and hotpotQA are plotted in Figure 3. The horizontal axis of each subplot depicts the compression ratio $\tau$ and the vertical axis corresponds to the metric score.
It can be observed that PartPrompt achieves considerable improvements for almost all compression ratios on these datasets, which again proves the superiority of the proposed method. Notably, the performance curve of PartPrompt displays a more pronounced convexity towards the upper left. This indicates that as the compression ratio decreases (i.e., the shorter compressed prompt), PartPrompt can better prioritize the deletion of less important parts, thereby slowing down the performance decline. It is worth mentioning the low to medium compression ratios are more challenging cases than the high compression ratios. And comparing to other methods, the proposed PartPrompt tends to receive more performance gains under the low and medium compression ratios. Also note that, the fluctuations on hotpotQA can be attributed to the brevity of the standard answers.
It is interesting to observe the similarity in the curves between BLEU and Rouge2, and between Rouge1 and RougeL. The similarity can be owing to the underlying principles of the metrics. BLEU is derived from the average of its 1-gram to 4-gram, which could be considered as 2.5-gram. This is similar to Rouge2, which also employs 2-gram, and both count only continuous matching words. Conversely, Rouge1 and RougeL account for non-continuous matching words in their computations and exhibit greater similarity to each other. The consistency of the underlying principles and results in turn renders the reliability of our experimental results.
@@ VI-C Comparative Study: Using Different LLMs for Inference
Mixtral-8x7B
20% token constraint
Selective-Context 3.58 31.63 9.22 28.73 60.64
LLMLingua 9.56 35.09 15.28 32.89 61.97
PartPrompt 13.69 43.54 22.87 41.05 67.58
30% token constraint
Selective-Context 6.68 37.49 13.54 34.15 64.32
LLMLingua 10.22 38.55 16.66 35.88 64.64
PartPrompt 18.38 48.57 28.39 46.21 70.20
50% token constraint
Selective-Context 14.26 46.74 22.63 43.41 69.53
LLMLingua 15.71 45.98 23.21 43.17 69.52
PartPrompt 28.28 56.48 38.30 54.29 74.57
@@ VI-D Ablation Study of PartPrompt
To elucidate the effect of each PartPrompt component, a comprehensive ablation study is carried out. A total of nine variants of PartPrompt are designed and evaluated, and their organization is illustrated by a tree in Figure 4.
@@
The complete PartPrompt is used as a point of departure for the sequential removal of components, so as to observe the effect of each component. Initially, the token value adjustment module is omitted, resulting in the hierarchical global tree degenerating into a global tree with a single virtual node. This modification is referred variant ①. Subsequently, the global parse tree is removed, keeping only the local parse trees, which is denoted as variant ②. In this instance, each local parse tree is pruned independently and then merged. Finally, the local parse trees are removed, with only information entropy used for prompt compression, which is denoted as variant ③. Note that, the complete PartPrompt and its variants ①, ②, and ③ all employ local approximated information entropy. When the information entropy approximation module is further removed, the complete PartPrompt and the corresponding variants ①, ②, ③ become ④, ⑤, ⑥, ⑦ respectively. These four variants can also be considered as an ablation chain of the global tree, and the performance difference among them can also reflect the contributions of each component of global tree.
@@
On the other hand, the variants ⑧ and ⑨ are developed to investigate the impact of entirely removing the information entropy module and solely utilizing the parse tree for compression. Regarding a compression ratio of 0.5 on BBCnews, the performance of the complete PartPrompt, all nine variants, as well as two baseline methods is shown in Table III.
@@
According to Table III, the complete PartPrompt achieves the best performance among all variants, demonstrating the effectiveness of each component. Notably, variants ⑧ and ⑨, which rely solely on the parse tree, also achieve commendable results. The variant ⑧ even outperforms two state-of-the-art selective methods: Selective-Context and LLMLingua. These observations imply that the parse tree, analyzed by linguistic rules, holds valuable information for prompt compression. Furthermore, the performance of variants ④, ⑤, ⑥, and ⑦ decreases as the corresponding component removes, which also reflects the positive contributions of each component.
@@
Figure 4 annotates the rise and drop of BS-F1 scores in the ablation study. First, it is evident that the token value adjustment yields a significant performance improvement (3.78%), which verifies the significance of considering human writing logic in prompt compression. Second, variants ⑧ and ⑨ consider the parse tree, while variants ① and ② consider both the parse tree and information entropy and thereby obtain better performance (0.35% and 1.73%) respectively. This observation indicates that the parse tree and the information entropy are complementary to each other in prompt compression.
Next, we provide an analysis for variants ④, ⑤, ⑥, ⑦ without information entropy approximation. In Section IV-G, we have concluded that the inter-sentence error of information entropy approximation should be managed or may result in performance loss. For example, the performance of variant ③ with approximation is inferior to that of variant 6 without approximation. Moreover, variant ② conducts the compression process within each sentence, which avoids the comparison of token values among inter-sentences and the inter-sentence error, thereby resulting in good performance. Unlike variant ②, variant ① employs the global tree that compares token values globally. In this case, the inter-sentence error would lead to the performance loss. Overall, variant ① can benefit from either not using the global tree (i.e., ②) or from not using the approximation (i.e., ⑤). Nonetheless, the complete PartPrompt do use the global tree as in variant ① and further exploits a node value adjustment mechanism to correct the approximation error, and receives the best performance.
@@
Figure 4: The organization of variants of PartPrompt in the ablation study. Each arrow comes with the removed component(s) and the corresponding performance drop (if ”--” or gain if ”$+$”). Each circled number ①-⑨ at the end of an arrow indicates the variant of PartPrompt after removing the component of that arrow based on the variant at the beginning of that arrow. The detailed performance of these variants can be found in Table III.
@@ VI-F Non-discourse and Discourse Metrics
To analyze whether the compressed prompt can effectively preserve the original prompt while maintaining coherence, a direct comparison is made between the compressed prompt and the original prompt, with both non-discourse and discourse metrics employed for evaluation. Non-discourse metrics is employed to measure the similarity of texts at the lexical level, including BLEU, Rouge, and BERTScore. Rouge (including Rouge1, Rouge2, and RougeL) measures the overlap between the hypothesis and reference texts. BERTScore (BS-F1 is used) offers a more comprehensive analysis of lexical similarity. BLEU consists of 1-gram, 2-gram, 3-gram, and 4-gram, where the n-gram calculates the proportion of n consecutive identical words. BLEU is employed to evaluate the continuity of the compressed text at the lexical level while measuring the lexical similarity. All the above non-discourse metrics are converted into percentages.
In contrast to non-discourse metrics, discourse metrics concern more on the global structure of texts, which can further evaluate the coherence of the compressed prompt [43, 44, 45]. Concretely, RC and LC [46] measure the number and proportion of words that serve the connecting role. EntityGraph [47] evaluates text coherence through graphs. LexicalChain [48] measures the overlap of lexical chains between the hypothesis and reference texts. DiscoScore (including DS-Focus and DS-SENT) measures the difference between the hypothesis and reference texts through their focus. We adopt the same settings as in [45] for these non-discourse metrics.
@@ VII Conclusion
In this work, we introduced a novel prompt compression method called PartPrompt, which leverages parse trees to guide the prompt compressing process. To this end, the prompt compression problem was transformed into a tree pruning problem. And during the prompt compression process, the linguistic rules and human writing logic were achieved via constructing the global parse tree and adjusting the node value in the tree. The comparative experiments demonstrated the state-of-the-art performance of PartPrompt across a wide range of datasets, compression ratios, metrics, and target LLMs for inference. A comprehensive ablation study verified the usefulness of each component in PartPrompt. And additional experiments further confirmed the advantages of PartPrompt in terms of more scenarios and metrics such as the extreme long prompt case and the coherence of compressed prompts.

### v3

@@ VI-B Comparative Study: Under Various Compression Ratios
The compression ratio is a vital aspect in the prompt compression problem. To further investigate the effectiveness of these methods across a range of compression ratios, we carry out the experiments with nine different compression ratios, i.e., $\tau=0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9$. The results for BBCnews, truncate arXiv, PeopleDaily and CodeNet are plotted in Figure 3. The horizontal axis of each subplot depicts the compression ratio $\tau$ and the vertical axis corresponds to the metric score.
It can be observed that PartPrompt achieves considerable improvements for almost all compression ratios on these datasets, which again proves the superiority of the proposed method. Notably, the performance curve of PartPrompt displays a more pronounced convexity towards the upper left. This indicates that as the compression ratio decreases (i.e., the shorter compressed prompt), PartPrompt can better prioritize the deletion of less important parts, thereby slowing down the performance decline. It is worth mentioning the low to medium compression ratios are more challenging cases than the high compression ratios. And comparing to other methods, the proposed PartPrompt tends to receive more performance gains under the low and medium compression ratios in BBCnews, truncate arXiv and PeopleDaily datasets. Furthermore, both Selective-Context and LLMLingua2 methods exhibit poor adaptability to PeopleDaily, the Chinese dataset, producing outputs with substantial unreadable texts. It leads to an actual compression ratio that is significantly greater than the target ratio and a significant decline in performance.
An interesting observation is the similarity in curves between BLEU and ROUGE-2, and between ROUGE-1 and ROUGE-L. This pattern alignment can be attributed to their underlying computational principles. Specifically, both BLEU and ROUGE-2 rely on contiguous n-gram matching. Conversely, ROUGE-1 and ROUGE-L incorporate non-contiguous word matches and exhibit greater similarity to each other. The consistency of the underlying principles and results in turn renders the reliability of our experimental results.
@@ VI-D Ablation Study of PartPrompt
To elucidate the effect of each PartPrompt component, a comprehensive ablation study is carried out. A total of ten variants of PartPrompt are designed and evaluated, and their organization is illustrated by a tree in Figure 4.
@@
The complete PartPrompt is used as a point of departure for the sequential removal of components, so as to observe the effect of each component. Initially, the first node adjustment technique is ablated, indicating that hyper-parameter $a_{2}$ was constrained to 1, noted as variant ⓪. Then the whole value adjustment module is omitted, resulting in the hierarchical global tree degenerating into a global tree with a single virtual node. This modification is referred variant ①. Subsequently, the global parse tree is removed, keeping only the local parse trees, which is denoted as variant ②. In this instance, each local parse tree is pruned independently and then merged. Finally, the local parse trees are removed, with only information entropy used for prompt compression, which is denoted as variant ③. Note that, the complete PartPrompt and its variants ①, ②, and ③ all employ local approximated information entropy. When the information entropy approximation module is further removed, the complete PartPrompt and the corresponding variants ①, ②, ③ become ④, ⑤, ⑥, ⑦ respectively. These four variants can also be considered as an ablation chain of the global tree, and the performance difference among them can also reflect the contributions of each component of global tree.
@@
On the other hand, the variants ⑧ and ⑨ are developed to investigate the impact of entirely removing the information entropy module and solely utilizing the parse tree for compression. Regarding a compression ratio of 0.5 on BBCnews, the performance of the complete PartPrompt, all ten variants, as well as four baseline methods is shown in Table III.
@@
According to Table III, the complete PartPrompt achieves the best performance among all variants, demonstrating the effectiveness of each component. Notably, variants ⑧ and ⑨, which rely solely on the parse tree, also achieve commendable results. The variant ⑧ even outperforms three baseline selective methods: Selective-Context, LLMLingua and LongLLMLingua. These observations imply that the parse tree, analyzed by linguistic rules, holds valuable information for prompt compression. Furthermore, the performance of variants ④, ⑤, ⑥, and ⑦ decreases as the corresponding component removes, which also reflects their positive contributions.
@@
Figure 4 annotates the rise and drop of BS-F1 scores in the ablation study. First, it is evident that the first node adjustment (comparing to the variant ⓪) yields a significant performance improvement (3.98%), which verifies the significance of considering human writing logic in prompt compression. Second, variants ⑧ and ⑨ consider the parse tree, while variants ① and ② consider both the parse tree and information entropy and thereby obtain better performance (0.35% and 1.73%) respectively. This observation indicates that the parse tree and the information entropy are complementary to each other in prompt compression.
Section IV-G established that the information entropy approximation introduces an inter-sentence error, which can cause performance degradation if not managed. This error manifests clearly in our ablation results. For instance, variant ③ (which uses approximation) underperforms compared to variant ⑥ (which does not). Furthermore, variant ②’s design inherently mitigates this issue. By performing compression within individual sentences, it avoids comparing token values across sentences, thereby eliminating the source of the inter-sentence error entirely. This explains its strong performance. In contrast, variant ①’s reliance on a global parse tree necessitates cross-sentence token comparisons. Consequently, it is susceptible to the inter-sentence error introduced by the approximation, leading to the observed performance gap between variants ① and ②, and also between ① and ⑤ (the non-approximated counterpart of ①). The performance of variant ⓪ further indicates that root propagation and leaf propagation alone are insufficient to correct errors and must be used together with first node adjustment. Crucially, the complete PartPrompt architecture addresses this challenge. While it employs the global tree like variant ① (and thus faces potential inter-sentence error), it incorporates the node value adjustment mechanism specifically to correct the approximation bias, which achieving the best overall performance.
@@
Figure 4: The organization of variants of PartPrompt in the ablation study. Each arrow comes with the removed component(s) and the corresponding performance drop (if ”-” or gain if ”++”). Each circled number ⓪-⑨ at the end of an arrow indicates the variant of PartPrompt after removing the component of that arrow based on the variant at the beginning of that arrow. The detailed performance of these variants can be found in Table III.
@@ VI-F Non-discourse and Discourse Metrics
To analyze whether the compressed prompt can effectively preserve the original prompt while maintaining coherence, a direct comparison is made between the compressed prompt and the original prompt, with both non-discourse and discourse metrics employed for evaluation. Non-discourse metrics is employed to measure the similarity of texts at the lexical level, including BLEU, Rouge, and BERTScore. Concretely, Rouge (including Rouge1, Rouge2, and RougeL) measures the overlap between the hypothesis and reference texts. BERTScore (BS-F1 is used) offers a more comprehensive analysis of lexical similarity. BLEU consists of 1-gram, 2-gram, 3-gram, and 4-gram, where the n-gram calculates the proportion of n consecutive identical words. BLEU is employed to evaluate the continuity of the compressed text at the lexical level while measuring the lexical similarity. All the above non-discourse metrics are converted into percentages.
In contrast to non-discourse metrics, discourse metrics concern more on the global structure of texts, which can further evaluate the coherence of the compressed prompt [49, 50, 51]. Specifically, RC and LC [52] measure the number and proportion of words that serve the connecting role. EntityGraph [53] evaluates text coherence through graphs. LexicalChain [54] measures the overlap of lexical chains between the hypothesis and reference texts. DiscoScore (including DS-Focus and DS-SENT) measures the difference between the hypothesis and reference texts through their focus. We adopt the same settings as in [51] for these non-discourse metrics.
@@ VII Conclusion
In this work, we introduced a novel prompt compression method called PartPrompt, which leverages parse trees to guide the prompt compressing process. To this end, the prompt compression problem was transformed into a tree pruning problem. And during the prompt compression process, the linguistic rules and human writing logic were achieved via constructing the global parse tree and adjusting the node value in the tree. The comparative experiments demonstrated the state-of-the-art performance of PartPrompt across various datasets (including different prompt task types and text types), compression ratios, metrics, and target LLMs for inference. A comprehensive ablation study verified the usefulness of each component in PartPrompt. And additional experiments further confirmed the advantages of PartPrompt in terms of more scenarios and metrics such as the extreme long prompt case and the coherence of compressed prompts.
@@
Acknowledgements
We would like to sincerely thank the anonymous reviewers and editors for their valuable advice. This work was funded by the National Key Research and Development Program of China (grant no. 2022YFF1202400), and was also supported by the Henan Province Medical Science and Technology Research Plan Major Project jointly built by the province and the ministry (grant no. SBGJ202401001).
Supplementary Materials
to the Paper
”Parse Trees Guided LLM Prompt Compression”
Appendix A Detailed Complexity Analysis
We provide detailed explanations for some conclusions in Section F here. To keep the exposition clear, we restate the notation:
$m$: number of sentences in the input prompt $R$.
$n$: number of tokens in each sentence (all sentences are assumed to have the same length).
$r_{j,i}$: the i-th token of the j-th sentence.
$C(\cdot)$: length function, so $C(R)=mn$.
$K$: a constant related to the model size used for entropy computation. Since PartPrompt and the three baseline methods all rely on models with 7B parameters, the constant K can be considered identical for all four methods.
Selective-context uses a small model to directly compute the information entropy of each token in the input prompt. Since the computational cost of a transformer model is proportional to the square of the input text length, we have $\mathcal{O}(\text{Selective-Context})=\mathcal{O}(Km^{2}n^{2})$,
LLMLingua splits the input prompt $R$ into a list of sentences $R_{1},R_{2},\cdots,R_{m}]$. For the tokens in the j-th sentence, the information entropy of each token is computed using the following method:
That is, for sentences $R_{<j}$, their compressed versions are used.
Again, considering the quadratic complexity of the transformer, together with the fact that part of the KV cache can be reused, and additionally assuming a uniform compression ratio $\tau$ for every sentence, we obtain $\mathcal{O}(\text{LLMLingua})=\mathcal{O}(Kmn^{2}(1+(m-1)\tau+(m-2)\tau^{2}))$.
Next, we provide a detailed analysis of the computational complexity of Algorithm 2: Recursive Tree Compression (corresponding to Line 8 in Algorithm 3).
Lines 2-7 of the algorithm are used to merge two solution lists. Line 3 iterates over $l$, and line 4 in fact iterates over $z$. It follows that for every combination $(i,j)$, where $0\leq i\leq C(Q)-1$ and $0\leq j\leq C(P)-1$, the pair $(Q_{i},P_{j})$ is traversed exactly once. Thus, the complexity of the function MergeSolution is $\mathcal{O}(C(Q)\cdot C(P))$.
Next, we consider the complexity of lines 11 and 13 of the algorithm for an input node $v_{i}$. Line 11 iterates over all children of $v_{i}$. Suppose $v_{i}$ has $b$ child nodes, denoted as $v^{(i)}_{1},v^{(i)}_{2},...,v^{(i)}_{k},...,v^{(i)}_{b}$. Let the length of the subtree corresponding to child node $v^{(i)}_{k}$ be $C(V^{(i)}_{k})$. The length of the subtree rooted at $v_{i}$ is $C(V_{i})=\sum_{j=1}^{b}C(V^{(i)}_{j})+C(v_{i})$.
[…]
$\sum_{k=1}^{b}\left(C(V^{(i)}_{k})\cdot\sum_{j=1}^{k-1}C(V^{(i)}_{j})\right)$ (15)
$=$ $\sum_{k=1}^{b}\sum_{j=1}^{k-1}\left(C(V^{(i)}_{k})C(V^{(i)}_{j})\right)$ (16)
$=$ $\frac{1}{2}\left(\left(\sum_{k=1}^{b}C(V^{(i)}_{k})\right)^{2}-\sum_{k=1}^{b}\left(C(V^{(i)}_{k}\right)^{2}\right)$ (17)
$\leq$ $\frac{1}{2}\left(\left(\sum_{k=1}^{b}C(V^{(i)}_{k})\right)^{2}-\frac{1}{b}\left(\sum_{k=1}^{b}C(V^{(i)}_{k})\right)^{2}\right)$ (18)
$=$ $\left(\frac{b-1}{2b}\right)\left(C(V_{i})-C(v_{i})\right)^{2}$ (19)
$\leq$ $\left(\frac{b-1}{2b}\right)\left(C(V_{i})\right)^{2}$ (20)
Therefore, when $C(V_{i})$ is fixed, the computational cost incurred by Line 13 does not exceed $\left(\frac{b-1}{2b}\right)\left(C(V_{i})\right)^{2}$. Equality holds when $C(V^{(i)}_{j})=C(V^{(i)}_{k}),\forall j,k$ and $C(v_{i})=0$. Furthermore, the computational cost of Lines 9-10 and 14-20 is $\mathcal{O}(1)$ and $\mathcal{O}(C(V_{i}))$, which is negligible. Since the algorithm is recursive, the computation required by Line 12 is essentially performed within the ”Line 13” corresponding to another node. Consequently, the overall computational cost of the algorithm is primarily attributed to the ”Line 13” computation for each node in the given tree. This cost is maximized when the subtree lengths of all child nodes are equal. Therefore, the total computational cost of the algorithm is maximized when the given tree is a balanced tree.
[…]
$=$ $\sum_{k=1}^{l}\left(\frac{b-1}{2b}\right)\left(b^{k-1}\right)^{2}b^{l-k}$ (21)
$=$ $\frac{1}{2}\left(b^{2l-2}-b^{l-2}\right)$ (22)
$<$ $\frac{b^{2l-2}}{2}$ (23)
The total length of all nodes in the tree $C(V)=b^{l-1}$. It follows directly that:
$\mathcal{O}(RecursiveTreeCompression)$ (24)
$\leq\mathcal{O}\left(C(V)^{2}\right)=\mathcal{O}(m^{2}n^{2})$ (25)
Appendix B Experiments on hotpotQA and GSM8K dataset
To test the performance of the proposed method on more diverse prompts, the hotpotQA[37] and GSM8K[39] datasets are also employed for additional testing. The answers in the dataset itself are used as ground truth. For hotpotQA dataset, we follow [38] to compute precision, recall, and F1 in addtional, since the answers is quite brief. For GSM8K dataset, as the questions are math-related, we adopt exact matching to calculate the EM score obeying the same setting as in [13, 7]. The related hyper-parameters are set as follows: for hotpotQA, $a_{1}=3$, $a_{2}=20$; for GSM8K, $a_{1}=5$, $a_{2}=25$. The experimental results are shown in Table VI. It is evident that PartPrompt achieves remarkable improvements in almost compression ratios, and metrics in both datasets, which demonstrates the advantages of PartPrompt.
TABLE VI: The performance of PartPrompt in comparison with baseline methods on hotpotQA and GSM8K datasets.
methods hotpotQA GSM8K
BLEU Rouge1 Rouge2 RougeL BS F1 precsion recall F1 Tokens $1/\tau$ EM Tokens $1/\tau$
20% tokens constraint
30% tokens constraint
50% tokens constraint
Appendix C Experiments on long prompt dataset
Many advanced LLMs work with longer prompts, making it necessary to evaluate the proposed method’s performance on longer prompts. We select the niah subset of RULER [40] and LongBench [41] datasets as benchmarks, sampling 100 test questions from each. The answers in the dataset itself are used as the ground truth. Due to the inherent length limitations of Selective-Context and LLMLingua, we therefore only compare our method against LLMLingua2 and LLM generative compression. For RULER-niah, we additionally report the performance of uncompressed prompts; this is infeasible for LongBench due to its excessive length. The hyper-parameters are set to $a_{1}=1,a_{2}=1$ for RULER-niah and $a_{1}=1,a_{2}=10$ for LongBench.
The experimental results are shown in the Table VII. It can be observed that on the RULER-niah dataset, PartPrompt significantly outperforms both baseline methods. Notably, both selective compression methods surpass the performance of the original uncompressed prompt, whereas LLM generation performs poorly. In fact, ”niah” stands for ”Needle-in-a-Haystack”. This dataset is designed to test a model’s ability to retrieve critical information (the ”needle”) from long sequences cluttered with distracting text (the ”haystack”). Therefore, a good compression method that can remove the distracting text while preserving the critical information simplifies the task, thereby enhancing the downstream model’s score. The experiments demonstrate that PartPrompt excels at this. In contrast, the uncompressed prompt’s excessive noise inherently challenges the LLM’s retrieval capability, explaining the poor performance of LLM generation compression. On the LongBench dataset, the proposed method also outperforms the compared methods, demonstrating its superiority.
TABLE VII: Performance of PartPrompt and compared methods on two longer datasets. Each dataset lists the scores and the length of the compressed prompt.
method RULER LongBench
score length score length
LLM Generation 0.0 2760.8 26.0 1918.2
LLMLingua2 24.0 2776.3 20.7 1972.3
PartPrompt 69.0 2150.0 26.4 1751.0
Original prompt 21.0 19191.9 NA NA
Appendix D Experiments on entropy model
The PartPrompt method employs a small model to calculate the information entropy of each token. In previous experiments, this is achieved using the Llama2-7B model. Considering that the model is an important part of the method, we show here a replacement experiment in which DeepSeek-LLM-7B[48] is employed to calculate the information entropy. We select the BBCnews dataset as the example, and the rest of the settings are the same as in Section VI-B. The results of the experiment can be found in Figure 6.
Figure 6: The performance of PartPrompt with another information entropy calculate model, DeepSeek-LLM-7B.
According to Figure 6, the proposed PartPrompt achieves nearly identical performance while using different information entropy models, and both versions of PartPrompt consistently outperform the baselines. These observations demonstrate the robustness of our method in terms of using different information entropy computation models.